# Two Heads are Better than One: Robust Learning Meets Multi-branch Models

### v2

@@ Abstract
Index Terms:
Deep neural network, Adversarial training, Model robustness, Multi-branch network.
@@ I Introduction
Deep Neural Networks (DNNs) have got great research attention and success in various application scenarios [1, 2]. However, it is found that DNNs are vulnerable to adversarial examples [3]. Leaving the model to produce the wrong output as the attacker wishes poses a significant security risk. Therefore, the vulnerability of DNNs to adversarial attacks is a critical topic in many areas with safety-vital nature, such as autonomous driving [4], and medical diagnostics [5].
@@
Ever since Szegedy et al. [7] brought this problem to the forefront, many studies have sprung up on generating adversarial examples [8, 9, 10] and building defenses against these attacks [3, 11, 12]. Among many successful defenses, adversarial training became one of the mainstream and proved to be effective and practical. Madry et al. [11] first introduced projected gradient descent (PGD) to generate adversarial examples for training, achieving model robustness and invariance to input perturbations. As a standard for training adversarially robust models, a variety of modifications over their implementation have been proposed, such as Lipschitz regularization [13, 14], computationally efficient adversarial training [15, 16], utilization of unlabeled data [17, 18] and better trade-offs between robustness and natural generalization [19, 20].
[…]
@@ II-A Adversarial Attacks
Research on the robustness of DNNs can be mainly divided into adversarial attacks and adversarial defenses, respectively. Adversarial attacks focus on crafting and adding imperceptible perturbation to the original input which makes the model misclassify. Adversarial attack techniques can be grouped into white-box and black-box methods. Under the setting of the white-box attack, the full architecture and parameters of DNNs are accessible while black-box attacks are not. Additionally, attacks may be classified into two groups based on the sort of misclassification attackers aim to achieve: targeted attacks where the attacker aims to influence the misclassification label and untargeted attacks where misclassification occurs regardless of any particular label. Goodfellow, Shlens, and Szegedy [3] proposed the Fast Gradient Sign Method (FGSM) which generates adversarial examples. It was followed by R+FGSM [24], which adds a randomization step, and the Basic Iterative Method (BIM) [25], which takes multiple smaller gradient steps.
In this paper, we mainly focus on AutoAttack [6], which is widely recognized as a standard evaluation for model robustness combined with four complementary attacks, to test our method and make comparison with others.
@@ II-B Adversarial Training as a defense
On the opposite to the thriving development of adversarial attack methods, researchers also demonstrate great interest in creating various defense methods, which can be roughly divided into two different approaches, the certified defense, and the empirical defense. While the certified defense adopts optimization objective (loss functions or regularization terms) to provide some guarantee over robustness [26, 27, 28], the empirical defense utilizes adversarial samples during the training process.
@@
To solve the inner maximization problem, PGD [11] is a normally used and effective method to generate perturbations. PGD updates perturbations on the clean data with a small step size for a regulated number of steps. The idea of performing attacks iteratively makes it a more powerful attack method over the Fast Gradient Sign Method (FGSM) [3] which only iterates once. Then, the generated adversarial data would be projected back onto a $l_{p}$-norm ball with radius $\epsilon$, $l_{\infty}$ and $l_{2}$ in a normal case.
Besides from adversarial training, there are many promising non-certified defense methods that argue robustness, such as the defense distillation [29], the activation puring [30], the input transformation [31], and etc. However, many of these methods were then proved to essentially cause obfuscated gradients that give a false sense of security [32]. So that these defense methods can still be breached by reconstructing the gradient or by using stronger black-box attack methods. On the contrary, methods based on adversarial training such as PAT [11], TRADES [19] and MART [33] still yield good performance on many white-box and black-box attacks.
Recently, Gowal et al. [23] studied the effectiveness of multiple hyper-parameters and uncovered the sizable robust-accuracy gap between models trained solely on data from the original training set and those trained with additional data extracted from the “80 Million Tiny Images” dataset (80M-TI). Gowal et al. [34] followed this work by proposing that even the non-realistic data generated by generative models can also provide a considerable improvement in robustness. Rebuffi et al. [35] conducted extensive experiments and made a comparison between heuristics-driven augmentation techniques and generative models.
@@ III Proposed Method
Notation
$(x,y)\in\mathcal{X}$ is the natural dataset, and $x^{\prime}$ signifies perturbed examples. $f^{k}$ denotes any of the $K$ branches and $f_{i}^{k}$ especially stands for the i-th ‘block’ (consistent to the definition in Zagoruyko and Komodakis [36]) of branch $k$. $\delta$ indicates perturbations bounded in the $\epsilon$-ball, where $\epsilon$ represents the numerical limit of the perturbation.
III-A Multi-branch Neural Network
Randomness might be enough
Routing to one of $K$ branches randomly is a thoroughly non-differentiable operation, so that white-box attacks that are based on optimization will be completely invalidated. To be more precise, every iteration for generating perturbations can only target a specific branch of the network, not the entire. Predictably, if we use some differentiable approach for branch selection, the multi-branch network will still be very vulnerable, and the multi-branch design will then become redundancy.
III-B Branch-orthogonal Loss
III-C Optimization Formulation
Loss for inner maximization
Retrospective of the mini-max optimization problem in adversarial training, the inner maximization is for generating adversarial examples $x^{\prime}$ from clean examples $x$. Here, we use the commonly used the cross-entropy loss to find adversarial examples as follows:
where $\mathrm{CE}$ denotes the cross-entropy loss.
Loss for outer minimization
The outer minimization problem is to allow the model to correctly classify adversarial examples. To obtain a better trade-off between the model robustness and its accuracy on clean data, we use the loss design proposed in TRADES[19]. In addition to the branch-orthogonal loss we introduced for orthogonality learning, the loss is composed of two parts: the first encourages the natural error (on clean data) to be minimized by applying the cross-entropy loss between the prediction of clean example $f^{k}(x)$ and labels $y$, and the second acts like a regularization term optimizing the difference between the prediction of clean example $f(x)$ and that of adversarial example $f^{k}(x^{\prime})$. And there is a hyper-parameter balancing the importance of the two parts.
End-to-end optimization
The full version of loss that we use to optimize networks is presented here, combining all of the aforementioned loss terms. It consists of three distinct components, each of which serves a different purpose: the cross-entropy loss to fit our model to clean examples, the Kullback-Leibler loss to close the gap between prediction results for clean and perturbed examples, and our newly proposed Branch-orthogonal loss to make the solution space of each branch orthogonal.
$\mathcal{L}_{\mathrm{clean}}=\mathrm{CE}(f^{k}(x),y),$ (4)
$\mathcal{L}_{\mathrm{KL}}=\mathrm{KL}(f^{k}(x^{\prime}),f^{k}(x)),$ (5)
@@ III-C Optimization Formulation
0: Model $f$, data $x$, labels $y$, num of branches $K$, number of epochs $epochs$, PGD algorithm $\mathrm{PGD}()$
@@ IV-A Evaluation Setting
Three publicly available benchmark datasets are used. The CIFAR-10 dataset consists of 60,000 32x32 color images of 10 classes, with 6,000 images per class. There are 50,000 training images and 10,000 test images. The CIFAR-100 dataset is like CIFAR-10, except that it has 100 classes, each containing 600 images, of which 500 are training images and 100 are test images. The Street View House Number (SVHN) Dataset is derived from Google street view house numbers, each image contains a set of ’0-9’ Arabic numbers. The training set contains 73257 digits and the test set contains 26032 digits.
FGSM, PGD, and AutoAttack are utilized to evaluate our proposed method along with state-of-the-arts. Testing robustness against PGD is done in 40/100 iterations with the maximum $\epsilon=8/255$ and step sizes of 0.010.010.01 for each iteration. FGSM attack also uses the same 8/255 bound for perturbation. We leverage the implementation of AutoAttack from torchattacks repository [42].
Model training details
We use WideResNets as our base models: WRN-28-10, WRN-34-10 and WRN-70-16 [36], respectively. Each branch is trained for 50 epochs using stochastic gradient descent (SGD) as the optimizer, with an initial learning rate of 0.1 that decays to 0.01 at the 40th epoch. The hyper-parameters are configured as $\lambda_{1}=\lambda_{2}=3$ during the first 40 epochs, followed by $\lambda_{1}=\lambda_{2}=1$ for the remaining 10 epochs.
@@ IV-B Overall Performance
Compared to all methods that do not use additional datasets, we obtain a maximum robustness improvement of about +7.23% in CIFAR-10, +9.07% in CIFAR-100 and +1.00%(WRN-28-10 solely) in SVHN over the highest results proposed. Our method also surpasses those trained with extra datasets: improvement of +3.95% and +8.45% in CIFAR-10 and CIFAR-100, respectively. In a longitudinal comparison of our method, it is consistent with the intuition that using a larger base model leads to an overall performance improvement. This improvement, however, is more minor than that provided by our proposed method.
@@ IV-C Ablation Studies
To further validate effectiveness of proposed Branch-orthogonal loss, a loss term to encourage the solution space of each branch orthogonal, we retrain a model completely independent of the BORT model, through which we implement the transfer attacks on four branches separately. The experiment results along with more detailed numbers in Table IV. From the table, it is clear that perturbations from TRADES-0 have a comparable impact on branches 0 and 3, suggesting that the solution spaces of TRADES-0 and TRADES-3 without Branch-orthogonal loss for training are fairly similar. This similarity in the solution space will degrade the apparent boost brought by integrating multiple solution spaces. The finding further demonstrates that Branch-orthogonal loss can minimize transfer attacks across branches of BORT by making the solution space of each branch as orthogonal as feasible, whereas the solution space produced through TRADES training without Branch-orthogonal loss is not at all guaranteed.
@@ V Conclusion
In this paper, we present BORT, the first adversarial training framework leveraging a multi-branch neural network, which improves network representation through multiple orthogonal branches while keeping inference time and computational cost from increasing, greatly improving model robustness against adversarial attacks. Through comprehensive experiments, we first illustrate that our proposed method goes well beyond the current state-of-the-art by a noticeable gap without the need for additional training datasets. The efficiency of multi-branch structure and branch-orthogonal loss for enhancing overall robustness was then discussed. Our method achieve top-rank performance on RobustBench under AutoAttack.

### v3

@@ I Introduction
Deep Neural Networks (DNNs) have got great research attention and success in various application scenarios [1, 2, 3, 4, 5, 6, 7, 8]. However, it is found that DNNs are vulnerable to adversarial examples [9, 10, 11, 12, 13]. Leaving the model to produce the wrong output as the attacker wishes poses a significant security risk. Therefore, the vulnerability of DNNs to adversarial attacks is a critical topic in many areas with safety-vital nature, such as autonomous driving [14, 15, 16], and medical diagnostics [17, 18, 19].
@@
Ever since Szegedy et al. [21] brought this problem to the forefront, many studies have sprung up on generating adversarial examples [22, 23] and building defenses against these attacks [24, 25]. Among many successful defenses, adversarial training became one of the mainstream and proved to be effective and practical. Madry et al. [24] first introduced projected gradient descent (PGD) to generate adversarial examples for training, achieving model robustness and invariance to input perturbations. As a standard for training adversarially robust models, a variety of modifications over their implementation have been proposed, such as Lipschitz regularization [26, 27], computationally efficient adversarial training [28, 29], utilization of unlabeled data [30, 31] and better trade-offs between robustness and natural generalization [32, 33].
However, in recent years, the improvement in robustness of the model seems to be less effective and practical. To further advance the model robustness, Carmon et al. [30], Najafi et al. [34] and Zhai et al. [35] pioneered the use of additional data (under both supervised and semi-supervised settings). Gowal et al. [36] trained a model on CIFAR-10 against $\ell_{\infty}$ perturbations of size $\epsilon=8/255$ with a maximum enhancement of $8.73\%$ robust accuracy when using additional unlabeled data, uncovering the importance of additional datasets for robustness of the model obtained from adversarial training.
@@ II-A Adversarial Attacks
Research on DNN robustness primarily involves adversarial attacks and defenses. Adversarial attacks generate subtle perturbations to input data, causing models to misclassify. These attacks are categorized as white-box (with full model access) or black-box (without model access), and as targeted (forcing a specific misclassification) or untargeted (any misclassification). Goodfellow, Shlens, and Szegedy [9] proposed the Fast Gradient Sign Method (FGSM) which generates adversarial examples. It was followed by R+FGSM [37], which adds a randomization step, and the Basic Iterative Method (BIM) [38], which takes multiple smaller gradient steps. In this paper, we mainly focus on AutoAttack [20], a widely used robustness benchmark that combines four complementary attacks, to evaluate and compare our method.
@@ II-B Adversarial Training as a defense
In response to the rapid development of adversarial attacks, researchers have proposed various defense methods, mainly categorized as certified and empirical defenses. Certified defenses use optimized objectives to guarantee robustness [39, 40, 41], while empirical defenses train models with adversarial samples.
@@
PGD [24] is a widely used method for generating adversarial perturbations by iteratively updating inputs and projecting them onto an $l_{p}$-norm ball, making it more effective than single-step attacks like FGSM [9].
Beyond adversarial training, various non-certified defenses (e.g., distillation [42], activation pruning [43], input transformation [44]) have been proposed, but many suffer from obfuscated gradients[45] and can be bypassed by advanced attacks. In contrast, adversarial training methods such as PAT[24], TRADES [32], and MART [46] remain effective against both white-box and black-box attacks. Recent work by Gowal et al.[36] revealed a robustness gap between models trained with and without extra data like 80M-TI. Follow-up studies[47, 48] showed that even synthetic data from generative models and various augmentation strategies can further enhance robustness.
@@ III Proposed Method
Notation. $(x,y)\in\mathcal{X}$ is the natural dataset, and $x^{\prime}$ signifies perturbed examples. $f^{k}$ denotes any of the $K$ branches and $f_{i}^{k}$ especially stands for the i-th ‘block’ (consistent to the definition in Zagoruyko and Komodakis [49]) of branch $k$. $\delta$ indicates perturbations bounded in the $\epsilon$-ball, where $\epsilon$ represents the numerical limit of the perturbation.
@@
0: Model $f$, data $x$, labels $y$, num of branches $K$, number of epochs epochsepochs, PGD algorithm $\mathrm{PGD}()$
@@
III-A Multi-branch Neural Network
Randomness might be enough. Routing to one of $K$ branches randomly is a thoroughly non-differentiable operation, so that white-box attacks that are based on optimization will be completely invalidated. To be more precise, every iteration for generating perturbations can only target a specific branch of the network, not the entire. Predictably, if we use some differentiable approach for branch selection, the multi-branch network will still be very vulnerable, and the multi-branch design will then become redundancy.
III-B Branch-orthogonal Loss
III-C Optimization Formulation
Loss for inner maximization. Retrospective of the mini-max optimization problem in adversarial training, the inner maximization is for generating adversarial examples $x^{\prime}$ from clean examples $x$. Here, we use the commonly used the cross-entropy loss to find adversarial examples as follows:
where $\mathrm{CE}$ denotes the cross-entropy loss.
Loss for outer minimization. The outer minimization problem is to allow the model to correctly classify adversarial examples. To obtain a better trade-off between the model robustness and its accuracy on clean data, we use the loss design proposed in TRADES[32]. In addition to the branch-orthogonal loss we introduced for orthogonality learning, the loss is composed of two parts: the first encourages the natural error (on clean data) to be minimized by applying the cross-entropy loss between the prediction of clean example $f^{k}(x)$ and labels $y$, and the second acts like a regularization term optimizing the difference between the prediction of clean example $f(x)$ and that of adversarial example $f^{k}(x^{\prime})$. And there is a hyper-parameter balancing the importance of the two parts.
End-to-end optimization. The full version of loss that we use to optimize networks is presented here, combining all of the aforementioned loss terms. It consists of three distinct components, each of which serves a different purpose: the cross-entropy loss to fit our model to clean examples, the Kullback-Leibler loss to close the gap between prediction results for clean and perturbed examples, and our newly proposed Branch-orthogonal loss to make the solution space of each branch orthogonal.
$\mathcal{L}_{\mathrm{clean}}=\mathrm{CE}(f^{k}(x),y),$ (4)
$\mathcal{L}_{\mathrm{KL}}=\mathrm{KL}(f^{k}(x^{\prime}),f^{k}(x)),$ (5)
@@ IV-A Evaluation Setting
Datasets and adversarial attacks. Three publicly available benchmark datasets are used. The CIFAR-10 dataset consists of 60,000 32x32 color images of 10 classes, with 6,000 images per class. There are 50,000 training images and 10,000 test images. The CIFAR-100 dataset is like CIFAR-10, except that it has 100 classes, each containing 600 images, of which 500 are training images and 100 are test images. The Street View House Number (SVHN) Dataset is derived from Google street view house numbers, each image contains a set of ’0-9’ Arabic numbers. The training set contains 73257 digits and the test set contains 26032 digits.
FGSM, PGD, and AutoAttack are utilized to evaluate our proposed method along with state-of-the-arts. Testing robustness against PGD is done in 40/100 iterations with the maximum $\epsilon=8/255$ and step sizes of $0.01$ for each iteration. FGSM attack also uses the same 8/255 bound for perturbation. We leverage the implementation of AutoAttack from torchattacks repository [55].
Model training details. We use WideResNets as our base models: WRN-28-10, WRN-34-10 and WRN-70-16 [49], respectively. Each branch is trained for 50 epochs using stochastic gradient descent (SGD) as the optimizer, with an initial learning rate of 0.1 that decays to 0.01 at the 40th epoch. The hyper-parameters are configured as $\lambda_{1}=\lambda_{2}=3$ during the first 40 epochs, followed by $\lambda_{1}=\lambda_{2}=1$ for the remaining 10 epochs.
@@ IV-B Overall Performance
As shown in Table I, our method achieves maximum robustness improvements of +7.23% on CIFAR-10, +9.07% on CIFAR-100, and +1.00% (WRN-28-10 only) on SVHN over the best results among methods that do not use additional datasets. Our method also surpasses those trained with extra datasets: improvement of +3.95% and +8.45% in CIFAR-10 and CIFAR-100, respectively. In a longitudinal comparison of our method, it is consistent with the intuition that using a larger base model leads to an overall performance improvement. This improvement, however, is more minor than that provided by our proposed method.
@@ IV-C Ablation Studies
To further validate the effectiveness of the proposed Branch-orthogonal loss, a loss term to encourage the solution space of each branch orthogonal, we retrain a model completely independent of the BORT model, through which we implement the transfer attacks on four branches separately. The experiment results along with more detailed numbers in Table IV. From the table, it is clear that perturbations from TRADES-0 have a comparable impact on branches 0 and 3, suggesting that the solution spaces of TRADES-0 and TRADES-3 without Branch-orthogonal loss for training are fairly similar. This similarity in the solution space will degrade the apparent boost brought by integrating multiple solution spaces. The finding further demonstrates that Branch-orthogonal loss can minimize transfer attacks across branches of BORT by making the solution space of each branch as orthogonal as feasible, whereas the solution space produced through TRADES training without Branch-orthogonal loss is not at all guaranteed.
@@ V Conclusion
In this paper, we present BORT, the first adversarial training framework leveraging a multi-branch neural network, which improves network representation through multiple orthogonal branches while keeping inference time and computational cost from increasing, greatly improving model robustness against adversarial attacks. Through comprehensive experiments, we first illustrate that our proposed method goes well beyond the current state-of-the-art by a noticeable gap without the need for additional training datasets. The efficiency of multi-branch structure and branch-orthogonal loss for enhancing overall robustness was then discussed. Our method achieves top-rank performance on RobustBench under AutoAttack.